The lengths of text messages are normally distributed with a population standard deviation of 3 characters and an unknown population mean. If a random sample of 27 text messages is taken and results in a sample mean of 22 characters, find a 98% confidence interval for the population mean. Round your answers to two decimal places. z0.10 z0.05 z0.04 z0.025 z0.01 z0.005 1.282 1.645 1.751 1.960 2.326 2.576 You may use a calculator or the common z-values above.
Answer #1

Solutions - Given - 0=3 x = 22 and n=27 Now, For 98% confidence interval C = 98% = 0.98 . =1-C = 2-0.98€ 0.02 4/2 = 0.02/2 =